OpenStudy (anonymous):

Help in need of desperate help A park is putting in a sidewalk of uniform width to go around two sides of a rectangular garden that is 10 feet by 30 feet. The contractor has enough concrete for 176 ft2. What is the maximum width of such a sidewalk?

OpenStudy (victoriasushchik):

h(t)=-16t^2+h , height is 50ft I need to know how to figure out how long it would take to fall to the ground When it hits the ground its height is zero. Solve:: -16t^2 + 50 = 0 16t^2 = 50 t^2 = 50/16 time = 1.77 seconds

OpenStudy (victoriasushchik):

@madmoiselle_sarcasm does this make sence?

OpenStudy (anonymous):

this is an area problem mor elike it actually I'm try to find "x" which is the width of the sidewalk

OpenStudy (victoriasushchik):

ohh ok

OpenStudy (anonymous):

yea sadly I need to do factoring as well in this too I just don't know how to exactly